OpenStudy (anonymous):

x^2 +xy + y^2 = 3 ,(-1,-1) tangent line to the curve

OpenStudy (anonymous):

differentiate it and find dy/dx....you got a slope and y-y1= m(x-x1

OpenStudy (anonymous):

find (dy/dx) by differentiating and then find the slope by putting the given points(-1,-1). equation of aline is gven by---> (y-y1)=m(x-x1) m=(dy/dx)=slope of the given line or tangent to a curve and u have the coordinates of a point on the line (x1,y1)=(-1,-1)